# Inference for the stochastic FitzHugh-Nagumo model from real action potential data via approximate Bayesian computation

### v1

@@ 3.2.2 Structure-based data summaries
The summary statistics employed here (cf. Lines 6 and 20 of Algorithm 1) are the structure-based summaries proposed in [6] for geometrically ergodic SDEs. In particular, we consider the invariant density $f_{\theta}$, and the invariant spectral density $S_{\theta}$ given by
@@
which is the Fourier transform of the auto-correlation function $R_{\theta}(\tau)$ with respect to the frequency $\nu$. Here, $f_{\theta}$ and $S_{\theta}$ are estimated based on a given dataset $\tilde{y}_{\theta}$, via a standard kernel density estimator $\hat{f}_{\tilde{y}_{\theta}}$ [48], and a smoothed periodogram estimator $\hat{S}_{\tilde{y}_{\theta}}$ [49], respectively, yielding the summaries
@@
Thanks to the ergodicity of the SDE, and its preservation by the numerical splitting scheme in Algorithm 2, these summaries have several desired features. First, they can be well estimated from a single path $\tilde{y}_{\theta}$ generated over a sufficiently long time horizon, instead of requiring a large number of repeated path simulations. Second, they are negligibly impacted by the (possibly unknown) initial value $X_{0}$. Moreover, as they are deterministic functions of the unknown parameter $\theta$, they fully characterise the asymptotic behaviour of the SDE (third), they are sensitive to small changes in $\theta$ (fourth), and robust to repeated simulations under the same parameter values (fifth). These characteristics are illustrated in Figure 1, where the top panel reports two paths obtained under the same value of $\theta$ (green and black trajectories), and a third path obtained under a slightly different value for $\theta$ (red trajectory). The corresponding summaries, derived from these datasets, are shown in the bottom panels. While judging on the similarities/differences between the trajectories seems difficult, this becomes straightforward when looking at their summaries, which are indistinguishable only when estimated from data generated under the same parameter value.
@@
A common alternative to the proposed summary statistics are what we call here canonical summaries, i.e., the mean, variance, skewness, kurtosis and auto-correlation up to a certain lag (here 5) of the original data and their differences, the latter to account for temporal dependence. This leads to a 18-dimensional summary vector which we denote by $s_{1}(\tilde{y}_{\theta})$ for a dataset $\tilde{y}_{\theta}$, and will only be considered in Section 4.2.
@@ 3.2.3 Choice of the distance measure
As a distance measure between the summaries of the observed dataset $y$ and synthetic dataset $\tilde{y}_{\theta}$ (cf. Lines 7 and 21 of Algorithm 1), we consider
@@
The Wasserstein distance between $y$ and $\tilde{y}_{\theta}$, as proposed in [2], may be used instead of $\textrm{IAE}(\hat{f}_{y},\hat{f}_{\tilde{y}_{\theta}})$, leading to marginal differences in the ABC posterior results (figures not shown). Moreover, as distance associated with the canonical summary statistics $s_{1}(\cdot)$, we consider the weighted Euclidean distance $d_{1}(s_{1}(y),s_{1}(\tilde{y}_{\theta}))=||\left(s_{1}(y)-s_{1}(\tilde{y}_{\theta})\right)/\alpha_{1}||$, where $\alpha_{1}\geq 0$ denotes a weight computed as mean absolute deviation of all summaries obtained during the first iteration of Algorithm 1 (i.e., the acceptance-rejection step).
@@ 3.2.4 Choice of prior distribution
The prior distribution is relevant in the acceptance-rejection step of Algorithm 1 ($r=1$, see Line 4), and later on, when updating the weights at the $r$th iteration (cf. Line 24). Here, we use two sets of uniform priors
@@
with the former used for the simulation study (in Section 4), and the latter for the real data study (in Section 5). Sampling $\gamma$ from a uniform in $(\epsilon/4,\cdot)$ guarantees that the condition $\kappa>0$ is met, see Remark 1.
@@ 3.2.5 Proposal Sampler
At the $r$th iteration, $\theta_{j}$, sampled from the weighted set $\{(\theta_{r-1}^{(j)},\omega_{r-1}^{(j)})_{j=1}^{N}\}$, is perturbed via a proposal sampler $K_{r}(\cdot|\theta_{j})$ to obtain $\theta_{j}^{*}$ (cf. Lines 17 and 18 of Algorithm 1.) Following [15], we choose $K_{r}(\cdot|\theta_{j})$ to be a multivariate normal perturbation kernel with mean $\theta_{j}$ and covariance matrix given by twice the empirical weighted covariance matrix $\hat{\Sigma}_{r-1}$ of the population $\{(\theta_{r-1}^{(j)},\omega_{r-1}^{(j)})_{j=1}^{N}\}$, i.e. $K_{r}(\cdot|\theta_{j})=\mathcal{N}(\theta_{j},2\hat{\Sigma}_{r-1})$. Hence, $k_{r}(\cdot|\theta_{r-1}^{(l)})$ in Line 24 of Algorithm 1 is then the density of $\mathcal{N}(\theta_{r-1}^{(l)},2\hat{\Sigma}_{r-1})$ evaluated in $\theta_{r}^{(j)},$ for $j,l\in\{1,\ldots,N\}$. This Gaussian proposal sampler, which we call here standard, is obviously not the only possible choice. In Appendix A, we also consider the optimal local covariance matrix (olcm) Gaussian proposal sampler introduced in [15], and compare the impact of the two samplers on the estimation results.
@@ 3.3 Implementation details
All algorithms are implemented in the software environment R [50], using the package Rcpp [13], offering a seamless integration of R and C++. In particular, Algorithm 2 is coded in C++, which speeds up the synthetic data generation significantly, reducing thus the computational cost of the ABC procedure. Taking advantage of the independence between the $N$ particles in the for loops in Algorithm 1, the code is parallelised using the packages foreach and doParallel. The R-package mvnfast [14] is used to draw samples from the multivariate normal perturbation kernel (using the function rmvn), and to compute its density (using the function dmvn).
The invariant spectral density $S_{\tilde{y}_{\theta}}$ and invariant density $f_{\tilde{y}_{\theta}}$ for a dataset $\tilde{y}_{\theta}$, used as summaries in (7), are estimated via the R-functions spectrum and density, respectively. Within the function spectrum, we avoid a logarithmic scale by setting log=“no”. Within the function density, we use the default value for the bandwidth bw, and set the number of points at which the density is estimated to n$=10^{3}$.
The number of particles accepted in each iteration is fixed to $N=10^{3}$ in all experiments. The threshold levels are chosen as the $50$th percentile (median) of the simulated (for the pilot study) or accepted (for $r\geq 1$) distances. All synthetic datasets (cf. Lines 5 and 19 of Algorithm 1) are generated with Algorithm 2, using the initial value $X_{0}=(0,0)^{\top}$, and the step-size $\Delta=0.02$.
An easy-to-grasp implementation of the proposed SBP SMC-ABC method (i.e., Algorithm 1 with the ingredients presented in Section 3.2, including Algorithm 2 for synthetic data simulation) is available at https://github.com/IreneTubikanec/SMC-ABC_FHN. Moreover, a comprehensive R-package, providing the codes for running different sequential ABC schemes, with different choices of the summary statistics, distances and proposal samplers (see Section 4.2 and the appendix) will be available upon publication.
All case studies, based on both simulated and real data, have been run on multiple core High-Performance-Clusters at both the University of Klagenfurt, and the University of Warwick.
@@ 4 Inference of the FHN parameters from simulated data
In this section, we investigate the performance of the proposed statistical procedure for parameter estimation in the stochastic FHN model (1). In Section 4.1, we analyse the SBP SMC-ABC method, based on different simulated observed datasets. In Section 4.2, we study the impact of summary statistics on the ABC results, illustrating how the structure-based summaries may outperform the canonical ones. Moreover, in Appendix A, we also extensively investigate the impact of proposal samplers in terms of acceptance rate, ESS, quality of the inference and running times. Other ingredients in the procedure, such as the percentile in defining the threshold levels (in particular, 10, 25, 50%), and alternative distances between the summaries (namely, the Wasserstein distance between the invariant densities instead of the IAE) have shown to have a limited impact on the results, and are therefore not reported here.
@@
Figure 2: Top panel: Observed (simulated) dataset for $\Delta_{\textrm{obs}}=0.02$ and $T_{\textrm{obs}}=50$ (yielding $n=2500$ data points, up to the blue line), $T_{\textrm{obs}}=100$ ($n=5000$, up to the red line), and $T_{\textrm{obs}}=200$ ($n=10^{4})$. Bottom panel: Observed (simulated) dataset for $\Delta_{\textrm{obs}}=0.08$ and $T_{\textrm{obs}}=50$ (yielding $n=625$ data points).
@@ 4.1 Inference under different observation settings
The performance of the proposed SBP SMC-ABC method is investigated under different observation regimes, depending on the choice of the observation time horizon $T_{\textrm{obs}}$, and the observation time step $\Delta_{\textrm{obs}}$. In particular, we consider $T_{\textrm{obs}}=50,100,200$, and $\Delta_{\textrm{obs}}=0.02,0.04,0.08$, yielding observed datasets with a size ranging from $n=10^{4}$ (for $T_{\textrm{obs}}=200$ and $\Delta_{\textrm{obs}}=0.02$) to only $n=625$ (for $T_{\textrm{obs}}=50$ and $\Delta_{\textrm{obs}}=0.08$) data points.
All observed datasets are obtained via Algorithm 2, using a simulation step-size $\Delta=10^{-4}$ and time horizon $T=200$. They are then sub-sampled with the corresponding observation time step $\Delta_{\textrm{obs}}$, and cut to the respective observation length $T_{\textrm{obs}}$, as visualised in Figure 2. The synthetic datasets are instead obtained with $T=T_{\textrm{obs}}$ and $\Delta=0.02$, and then sub-sampled with the corresponding $\Delta_{\textrm{obs}}$. In this section, the computational budget of Algorithm 1 is set to $\textrm{Nsim}_{\textrm{max}}=10^{6}$ synthetic data simulations. Moreover, the results reported here are based on one trial per observation setting. However, we verified that they are consistent across repeated trials (figures not shown).
Figure 3: $90$% CIs of the marginal posterior densities of the parameter vector $\theta$ (2) of the stochastic FHN model (1) as a function of the computational budget. Top panels: $T_{\textrm{obs}}=200$, and $\Delta_{\textrm{obs}}=0.02$ (black squares), $0.04$ (red triangles), $0.08$ (blue circles). Bottom panels: $\Delta_{\textrm{obs}}=0.02$, and $T_{\textrm{obs}}=200$ (black squares), $100$ (red triangles), $50$ (blue circles). The green horizontal lines and the dashed lines indicate the true parameter values and the corresponding posterior medians, respectively.
@@
Our inferential results show that the observation time step $\Delta_{\textrm{obs}}$ does not significantly influence the performance of the SBP SMC-ABC method, when choosing a (large enough) observation time horizon $T_{\textrm{obs}}$. This becomes evident when looking at the top panels of Figure 3, which show the $90$% credible intervals (CIs) of the derived marginal ABC posterior densities of $\theta$ (2) as a function of the computational budget, when fixing $T_{\textrm{obs}}=200$, and considering $\Delta_{\textrm{obs}}=0.02$ (black squares), $\Delta_{\textrm{obs}}=0.04$ (red triangles) and $\Delta_{\textrm{obs}}=0.08$ (blue circles). This is also the case when choosing $T_{\textrm{obs}}=50$ or $T_{\textrm{obs}}=100$.
In contrast to that, the observation time horizon $T_{\textrm{obs}}$ does influence the inferential results when fixing the time-step $\Delta_{\textrm{obs}}$, with smaller $T_{\textrm{obs}}$ yielding broader CIs. This can be observed in the bottom panels of Figure 3, where we fix $\Delta_{\textrm{obs}}=0.02$, and consider $T_{\textrm{obs}}=200$ (black squares), $T_{\textrm{obs}}=100$ (red triangles), and $T_{\textrm{obs}}=50$ (blue circles). Similar considerations hold when choosing $\Delta_{\textrm{obs}}=0.04$, or $\Delta_{\textrm{obs}}=0.08$.
It is important to stress that the number of observed data points $n$ is the same (for the same colours and shapes) in the top and bottom panels of Figure 3, i.e., $n=2500$ (blue circles), $n=5000$ (red triangles) and $n=10000$ (black squares). However, the increased sample size leads to a more accurate ABC inference (in the sense of narrower CIs) only when increasing the time horizon. This is also confirmed by Table 1, where the ABC posterior means and standard deviations obtained under the different observation settings are reported. These founding are in agreement with the requirement $T_{\textrm{obs}}\to\infty$ for contrast estimators of SDEs to be consistent and asymptotically normal (see, e.g., [9]).
@@
In Figure 4, we report the marginal ABC posterior densities of $\theta$ (2), derived from the two “uttermost” observation settings: (i) $T_{\textrm{obs}}=200$, $\Delta_{\textrm{obs}}=0.02$, $n=10^{4}$ (black solid lines); (ii) $T_{\textrm{obs}}=50$, $\Delta_{\textrm{obs}}=0.08$, $n=625$ (red dashed lines). In both scenarios, the marginal posterior densities show a clear update compared to the respective uniform prior densities (blue horizontal lines), and cover the true parameter values (green vertical lines), with the posteriors of scenario (i) being narrower than those of scenario (ii) (cf. the values for the corresponding ABC posterior standard deviations in Table 1).
Table 1: ABC posterior means and standard deviations of $\theta$ (2), under different observation settings. The true parameter values are $\epsilon=0.1$, $\gamma=1.5$, $\beta=0.8$, and $\sigma=0.3$.
$T_{\textrm{obs}}$ $\Delta_{\textrm{obs}}$ $n$ ABC posterior means ABC posterior stand. dev.s
$200$ $0.02$ $10^{4}$ $(0.101,1.523,0.808,0.300)$ $(0.010,0.087,0.062,0.023)$
$200$ $0.04$ $5000$ $(0.101,1.524,0.808,0.300)$ $(0.009,0.086,0.061,0.022)$
$200$ $0.08$ $2500$ $(0.102,1.531,0.812,0.301)$ $(0.010,0.089,0.064,0.023)$
$100$ $0.02$ $5000$ $(0.101,1.504,0.787,0.304)$ $(0.014,0.119,0.085,0.034)$
$100$ $0.04$ $2500$ $(0.102,1.508,0.789,0.305)$ $(0.014,0.125,0.089,0.033)$
$100$ $0.08$ $1250$ $(0.101,1.506,0.787,0.304)$ $(0.013,0.116,0.084,0.033)$
$50$ $0.02$ $2500$ $(0.096,1.480,0.787,0.284)$ $(0.017,0.160,0.113,0.040)$
$50$ $0.04$ $1250$ $(0.098,1.486,0.789,0.281)$ $(0.018,0.166,0.125,0.039)$
$50$ $0.08$ $625$ $(0.098,1.490,0.793,0.282)$ $(0.018,0.171,0.123,0.041)$

### v2

@@ 3.2.2 Structure-based data summaries
The summary statistics employed here (cf. Lines 6 and 21 of Algorithm 1) are the structure-based summaries proposed in [6] for geometrically ergodic SDEs. In particular, we consider the invariant density $f_{\theta}$ and the invariant spectral density $S_{\theta}$ given by
@@
which is the Fourier transform of the auto-correlation function $R_{\theta}(\tau)$ with respect to the frequency $\nu$. Here, $f_{\theta}$ and $S_{\theta}$ are estimated based on a given dataset $\tilde{y}_{\theta}$, via a standard kernel density estimator $\hat{f}_{\tilde{y}_{\theta}}$ [51] and a smoothed periodogram estimator $\hat{S}_{\tilde{y}_{\theta}}$ [53], respectively, yielding the summaries
@@
Thanks to the ergodicity of the SDE and its preservation by the numerical splitting scheme in Algorithm 2, these summaries have several desired features. First, they can be well estimated from a single path $\tilde{y}_{\theta}$ generated over a sufficiently long time horizon, instead of requiring a large number of repeated path simulations. Second, they are negligibly impacted by the (possibly unknown) initial value $X_{0}$. Moreover, as they are deterministic functions of the unknown parameter $\theta$, they fully characterise the asymptotic behaviour of the SDE (third), they are sensitive to small changes in $\theta$ (fourth) and robust to repeated simulations under the same parameter values (fifth). These characteristics are illustrated in Figure 1, where the top panel reports two paths obtained under the same value of $\theta$ (green and black trajectories) and a third path obtained under a slightly different value for $\theta$ (red trajectory). The corresponding summaries, derived from these datasets, are shown in the bottom panels. While judging on the similarities/differences between the trajectories seems difficult, this becomes straightforward when looking at their summaries, which are indistinguishable only when estimated from data generated under the same parameter value.
@@
A common alternative to the proposed summary statistics are what we call here canonical summaries, i.e., the mean, variance, skewness, kurtosis and auto-correlation up to a certain lag (here 5) of the original data and their differences, the latter to account for temporal dependence. This leads to a 18-dimensional summary vector which we denote by $s_{1}(\tilde{y}_{\theta})$ for a dataset $\tilde{y}_{\theta}$, and will only be considered in Section 4.2 and Appendix A.
@@ 3.2.3 Choice of the distance measure
As a distance measure between the summaries of the observed dataset $y$ and synthetic dataset $\tilde{y}_{\theta}$ (cf. Lines 7 and 22 of Algorithm 1), we consider
@@
The Wasserstein distance between $y$ and $\tilde{y}_{\theta}$, as proposed in [2], may be used instead of $\textrm{IAE}(\hat{f}_{y},\hat{f}_{\tilde{y}_{\theta}})$, leading to marginal differences in the ABC posterior results (figures not shown). Moreover, as distance associated with the canonical summary statistics $s_{1}(\cdot)$, we consider the weighted Euclidean distance $d_{1}(s_{1}(y),s_{1}(\tilde{y}_{\theta}))=||\left(s_{1}(y)-s_{1}(\tilde{y}_{\theta})\right)/\alpha_{1}||$, where $\alpha_{1}$ denotes a vector of weights with $\min(\alpha_{1})>0$, the division by $\alpha_{1}$ being interpreted component-wise [52]. The vector $\alpha_{1}$ is computed as mean absolute deviations of all summaries obtained during the first iteration of Algorithm 1 (i.e., the acceptance-rejection step), as done, e.g., in [49]. Other possible weights can be chosen (e.g. median absolute deviation [8]), all motivated by the idea of normalising the effect caused by having summary statistics varying on very different scales [52].
@@ 3.2.4 Choice of prior distribution
The prior distribution is relevant in the acceptance-rejection step of Algorithm 1 ($r=1$, see Line 4), and later on, when updating the weights at the $r$th iteration (cf. Line 25). Here, we use two sets of uniform priors
@@
with the former used for the simulation study (in Section 4) and the latter for the real data study (in Section 5). Sampling $\gamma$ from a uniform in $(\epsilon/4,\cdot)$ guarantees that the condition $\kappa>0$ is met, see Remark 1.
The use of continuous uniform priors is inspired by [6]. However, other choices are possible. In Appendix B, we also consider log-normal and exponential priors, and illustrate that their impact on the estimation results is marginal.
@@ 3.2.5 Proposal Sampler
At the $r$th iteration, $\theta_{j}$, sampled from the weighted set $\{(\theta_{r-1}^{(j)},\omega_{r-1}^{(j)})_{j=1}^{N}\}$, is perturbed via a proposal sampler $K_{r}(\cdot|\theta_{j})$ to obtain $\theta_{j}^{*}$ (cf. Lines 17 and 18 of Algorithm 1). Following [17], we choose $K_{r}(\cdot|\theta_{j})$ to be a multivariate normal perturbation kernel with mean $\theta_{j}$ and covariance matrix given by twice the empirical weighted covariance matrix $\hat{\Sigma}_{r-1}$ of the population $\{(\theta_{r-1}^{(j)},\omega_{r-1}^{(j)})_{j=1}^{N}\}$, i.e. $K_{r}(\cdot|\theta_{j})=\mathcal{N}(\theta_{j},2\hat{\Sigma}_{r-1})$. Hence, $k_{r}(\cdot|\theta_{r-1}^{(l)})$ in Line 25 of Algorithm 1 is then the density of $\mathcal{N}(\theta_{r-1}^{(l)},2\hat{\Sigma}_{r-1})$ evaluated in $\theta_{r}^{(j)},$ for $j,l\in\{1,\ldots,N\}$.
This Gaussian proposal sampler, which we call here standard, is obviously not the only possible choice. In Appendix C, we also consider the optimal local covariance matrix (olcm) Gaussian proposal sampler introduced in [17], and compare the impact of the two samplers on the estimation results.
Differently than from the prior, the condition $\kappa>0$ is not directly accounted for in the proposal samplers $K_{r}$. However, perturbed samples $\theta^{*}$ which do not satisfy $\kappa>0$ or lie outside the support of the prior (i.e. $\pi(\theta^{*})=0$) are immediately rejected and replaced by new ones (cf. Line 19 of Algorithm 1 and the right panel of Figure 11 in Appendix A). This is a standard procedure in sequential ABC samplers, see e.g. Step 4 of the SMC-ABC Algorithm 1 in [58] as well as Line 18 of the Sequential Important Sampling ABC (SIS-ABC) Algorithm 1 and Line 19 of the SMC-ABC Algorithm 2 in [49]. Immediate rejections due to entries of $\theta^{*}$ exceeding the upper bound of the respective uniform prior’s support can be avoided by using priors with support $(0,\infty)$ instead (cf. Appendix B). Rejections due to entries of $\theta^{*}$ being negative may be avoided by adjusting the algorithm to account for logged parameters. Both adjustments have a marginal impact on the estimation results and computing times. Also the number of particles sampled from $K_{r}$ with $\kappa\leq 0$ and thus immediately disregarded is negligible.
@@ 3.3 Implementation details
All algorithms are implemented in the software environment R [54], using the package Rcpp [15], offering a seamless integration of R and C++. In particular, Algorithm 2 is coded in C++, which speeds up the synthetic data generation significantly, reducing thus the computational cost of the ABC procedure. Taking advantage of the independence between the $N$ particles in the for loops in Algorithm 1, the code is parallelised using the packages foreach and doParallel. The R-package mvnfast [16] is used to draw samples from the multivariate normal perturbation kernel (using the function rmvn) and to compute its density (using the function dmvn).
The invariant spectral density $S_{\tilde{y}_{\theta}}$ and invariant density $f_{\tilde{y}_{\theta}}$ for a dataset $\tilde{y}_{\theta}$, used as summaries in (8), are estimated via the R-functions spectrum and density, respectively. Within the function spectrum, we avoid a logarithmic scale by setting log=“no”. Within the function density, we use the default value for the bandwidth bw and set the number of points at which the density is estimated to n$=10^{3}$.
The number of particles accepted in each iteration is fixed to $N=10^{3}$ in all experiments. The threshold levels are chosen as the $50$th percentile (median) of the simulated (for the pilot study) or accepted (for $r\geq 1$) distances. All synthetic datasets (cf. Lines 5 and 20 of Algorithm 1) are generated with Algorithm 2, using the step-size $\Delta=0.02$.
An easy-to-grasp implementation of the SBP SMC-ABC method (i.e., Algorithm 1 with the ingredients presented in Section 3.2, including Algorithm 2 for synthetic data simulation) is available at https://github.com/IreneTubikanec/SMC-ABC_FHN. Moreover, a comprehensive R-package, providing the codes for running Algorithm 1 with different summary statistics (structure-based and canonical), priors (uniform, log-normal and exponential), distances (IAE and Wasserstein) and proposal samplers (Gaussian samplers standard and olcm, described in Section 3.2.5 and Appendix C, respectively) is available at https://github.com/massimilianotamborrino/SMCABCFHN.
All case studies, based on both simulated and real data, have been run on multiple core High Performance Computing (HPC) clusters at both the University of Klagenfurt and the University of Warwick. Unless differently stated, all reported runtimes were obtained by launching the relevant codes on $48$ parallel cores of a HPC cluster at the University of Warwick (Dell PowerEdge C6420, 2.9 GHz, 24-core processors).
@@ 4 Inference of the FHN parameters from simulated data
In this section, we investigate the performance of the proposed statistical procedure for parameter estimation in the stochastic FHN model (1). In Section 4.1, we analyse the SBP SMC-ABC method, based on different simulated observed datasets. In Section 4.2, we study the impact of summary statistics on the ABC results, illustrating how the structure-based summaries may outperform the canonical ones. In Appendix A, we investigate the computational cost of the different summaries. In Appendix B, we illustrate the marginal impact of the prior distribution on the ABC results. Moreover, in Appendix C, we also extensively investigate the impact of proposal samplers in terms of acceptance rate, ESS, quality of the inference and running times. Other ingredients in the procedure, such as the percentile in defining the threshold levels (in particular, 10, 25, 50%) and alternative distances between the summaries (namely, the Wasserstein distance between the invariant densities instead of the IAE) have shown to have a limited impact on the results, and are therefore not reported here.
@@ 4.1 Inference under different observation settings
Figure 2: Top panel: Observed (simulated) dataset for $\Delta_{\textrm{obs}}=0.02$ and $T_{\textrm{obs}}=50$ (yielding $n=2500$ data points, up to the blue line), $T_{\textrm{obs}}=100$ ($n=5000$, up to the red line) and $T_{\textrm{obs}}=200$ ($n=10^{4})$. Bottom panel: Observed (simulated) dataset for $\Delta_{\textrm{obs}}=0.08$ and $T_{\textrm{obs}}=50$ (yielding $n=625$ data points).
The performance of the proposed SBP SMC-ABC method is investigated under different observation regimes, depending on the choice of the observation time horizon $T_{\textrm{obs}}$, and the observation time step $\Delta_{\textrm{obs}}$. In particular, we consider $T_{\textrm{obs}}=50,100,200$ and $\Delta_{\textrm{obs}}=0.02,0.04,0.08$, yielding observed datasets with a size ranging from $n=10^{4}$ (for $T_{\textrm{obs}}=200$ and $\Delta_{\textrm{obs}}=0.02$) to only $n=625$ (for $T_{\textrm{obs}}=50$ and $\Delta_{\textrm{obs}}=0.08$) data points.
All observed datasets are obtained via Algorithm 2, using a simulation step-size $\Delta=10^{-4}$ and time horizon $T=200$. They are then sub-sampled with the corresponding observation time step $\Delta_{\textrm{obs}}$ and cut to the respective observation length $T_{\textrm{obs}}$, as visualised in Figure 2. The synthetic datasets are instead obtained with $T=T_{\textrm{obs}}$ and $\Delta=0.02$, and then sub-sampled with the corresponding $\Delta_{\textrm{obs}}$. In this section, the computational budget of Algorithm 1 is set to $\textrm{Nsim}_{\textrm{max}}=10^{6}$ synthetic data simulations. Moreover, the results reported here are based on one trial per observation setting. However, we verified that they are consistent across repeated runs (figures not shown).
Figure 3: $90$% CIs of the marginal posterior densities of the parameter vector $\theta$ (2) of the stochastic FHN model (1) as a function of the computational budget. Top panels: $T_{\textrm{obs}}=200$ and $\Delta_{\textrm{obs}}=0.02$ (black squares), $0.04$ (red triangles), $0.08$ (blue circles). Bottom panels: $\Delta_{\textrm{obs}}=0.02$ and $T_{\textrm{obs}}=200$ (black squares), $100$ (red triangles), $50$ (blue circles). The green horizontal lines and the dashed lines indicate the true parameter values and the corresponding posterior medians, respectively. The number of observed data points is the same (for the same colours and shapes) for the top and bottom panels.
Table 1: ABC posterior means and standard deviations of $\theta$ (2), under different observation settings. The true parameter values are $\epsilon=0.1$, $\gamma=1.5$, $\beta=0.8$ and $\sigma=0.3$. The reported runtimes are medians across ten runs of the algorithm.
$T_{\textrm{obs}}$ $\Delta_{\textrm{obs}}$ $n$ ABC posterior means ABC posterior stand. dev.s Median runtime
$200$ $0.02$ $10^{4}$ $(0.101,1.523,0.808,0.300)$ $(0.010,0.087,0.062,0.023)$ 3 min 7 s
$200$ $0.04$ $5000$ $(0.101,1.524,0.808,0.300)$ $(0.009,0.086,0.061,0.022)$ 2 min 3 s
$200$ $0.08$ $2500$ $(0.102,1.531,0.812,0.301)$ $(0.010,0.089,0.064,0.023)$ 1 min 41 s
$100$ $0.02$ $5000$ $(0.101,1.504,0.787,0.304)$ $(0.014,0.119,0.085,0.034)$ 2 min 25 s
$100$ $0.04$ $2500$ $(0.102,1.508,0.789,0.305)$ $(0.014,0.125,0.089,0.033)$ 1 min 47s
$100$ $0.08$ $1250$ $(0.101,1.506,0.787,0.304)$ $(0.013,0.116,0.084,0.033)$ 1 min 26 s
$50$ $0.02$ $2500$ $(0.096,1.480,0.787,0.284)$ $(0.017,0.160,0.113,0.040)$ 1 min 25 s
$50$ $0.04$ $1250$ $(0.098,1.486,0.789,0.281)$ $(0.018,0.166,0.125,0.039)$ 1 min 8 s
$50$ $0.08$ $625$ $(0.098,1.490,0.793,0.282)$ $(0.018,0.171,0.123,0.041)$ 1 min 3 s
@@
Figure 5: Scatter plots of pairs of the kept ABC posterior samples. Black dots: $T_{\textrm{obs}}=200$, $\Delta_{\textrm{obs}}=0.02$, $n=10^{4}$. Red dots: $T_{\textrm{obs}}=50$, $\Delta_{\textrm{obs}}=0.08$, $n=625$. The green lines indicate the respective true parameter values.
Our inferential results show that the observation time step $\Delta_{\textrm{obs}}$ does not significantly influence the performance of the SBP SMC-ABC method, when choosing a (large enough) observation time horizon $T_{\textrm{obs}}$. This becomes evident when looking at the top panels of Figure 3, which show the $90$% credible intervals (CIs) of the derived marginal ABC posterior densities of $\theta$ (2) as a function of the computational budget, when fixing $T_{\textrm{obs}}=200$ and considering $\Delta_{\textrm{obs}}=0.02$ (black squares), $\Delta_{\textrm{obs}}=0.04$ (red triangles) and $\Delta_{\textrm{obs}}=0.08$ (blue circles). This is also the case when choosing $T_{\textrm{obs}}=50$ or $T_{\textrm{obs}}=100$.
In contrast to that, the observation time horizon $T_{\textrm{obs}}$ does influence the inferential results when fixing the time-step $\Delta_{\textrm{obs}}$, with smaller $T_{\textrm{obs}}$ yielding broader CIs. This can be observed in the bottom panels of Figure 3, where we fix $\Delta_{\textrm{obs}}=0.02$ and consider $T_{\textrm{obs}}=200$ (black squares), $T_{\textrm{obs}}=100$ (red triangles) and $T_{\textrm{obs}}=50$ (blue circles). Similar considerations hold when choosing $\Delta_{\textrm{obs}}=0.04$, or $\Delta_{\textrm{obs}}=0.08$.
It is important to stress that the number of observed data points $n$ is the same (for the same colours and shapes) in the top and bottom panels of Figure 3, i.e., $n=2500$ (blue circles), $n=5000$ (red triangles) and $n=10000$ (black squares). However, the increased sample size leads to a more accurate ABC inference (in the sense of narrower CIs) only when increasing the time horizon. This is also confirmed by Table 1, where the ABC posterior means and standard deviations obtained under the different observation settings are reported together with the median runtimes obtained over ten runs of the SBP SMC-ABC algorithm. These findings are in agreement with the requirement $T_{\textrm{obs}}\to\infty$ for contrast estimators of SDEs to be consistent and asymptotically normal (see, e.g., [10]).
@@
In Figure 4, we report the marginal ABC posterior densities of $\theta$ (2), derived from the two most extreme observation settings: (i) $T_{\textrm{obs}}=200$, $\Delta_{\textrm{obs}}=0.02$, $n=10^{4}$ (black solid lines); (ii) $T_{\textrm{obs}}=50$, $\Delta_{\textrm{obs}}=0.08$, $n=625$ (red dashed lines). In both scenarios, the marginal posterior densities show a clear update compared to the respective uniform prior densities (blue horizontal lines) and cover the true parameter values (green vertical lines), with the posteriors of scenario (i) being narrower than those of scenario (ii) (cf. the values for the corresponding ABC posterior standard deviations in Table 1).
To provide a more comprehensive picture of the obtained ABC posterior distributions, in Figure 5, we also report scatter plots of pairs of the kept ABC posterior samples for scenarios (i) (black dots) and (ii) (red dots), with similar results obtained in the other settings. The underlying true parameter values, which are represented by the green vertical and horizontal lines, are covered by the respective point clouds. All pairs of parameters show some correlation, with $\gamma$ and $\beta$ strongly correlated, having a correlation coefficient of $0.912$ and $0.926$ for scenario (i) and (ii), respectively.